OpenStudy (nuccioreggie):

Describe the distributive property? How is it different from other properties? Can you give an Example?

OpenStudy (nuccioreggie):

@Abhisar

OpenStudy (nuccioreggie):

@iGreen. @Compassionate

OpenStudy (nuccioreggie):

@jakashaka123

OpenStudy (timbo5lice):

The Distributive Property is easy to remember, if you recall that "multiplication distributes over addition". Formally, they write this property as "a(b + c) = ab + ac". In numbers, this means, that 2(3 + 4) = 2×3 + 2×4. Any time they refer in a problem to using the Distributive Property, they want you to take something through the parentheses (or factor something out); any time a computation depends on multiplying through a parentheses (or factoring something out), they want you to say that the computation used the Distributive Property. Why is the following true? 2(x + y) = 2x + 2y Since they distributed through the parentheses, this is true by the Distributive Property. Use the Distributive Property to rearrange: 4x – 8 The Distributive Property either takes something through a parentheses or else factors something out. Since there aren't any parentheses to go into, you must need to factor out of. Then the answer is "By the Distributive Property, 4x – 8 = 4(x – 2)" "But wait!" you say. "The Distributive Property says multiplication distributes over addition, not subtraction! What gives?" You make a good point. This is one of those times when it's best to be flexible. You can either view the contents of the parentheses as the subtraction of a positive number ("x – 2") or else as the addition of a negative number ("x + (–2)"). In the latter case, it's easy to see that the Distributive Property applies, because you're still adding; you're just adding a negative. The other two properties come in two versions each: one for addition and the other for multiplication. (Note that the Distributive Property refers to both addition and multiplication, too, but to both within just one rule.) HOPE THIS HELPS

OpenStudy (anonymous):

@nuccioreggie srrry bruh i was off line

OpenStudy (nuccioreggie):

Its alright man ima just tag in ok

OpenStudy (timbo5lice):

did i help any??

OpenStudy (nuccioreggie):

How is distributive property different from other Properties?